solve by elimination
2c - 3d = 14
5c - 6d = 32

Answer by stanbon(75887)

solve by elimination
2c - 3d = 14
5c - 6d = 32
-----------------------
Multiply thru the top equation by 2 to get:
4c - 6d = 28
5c - 6d = 32
------------------------
Subtract and solve for "c":
c = 4
---
Solve for "d":
2c-3d = 14
2*4 - 3d = 14
-3d = 6
d = -2
=================
Cheers,
Stan H.
=================